OpenStudy (anonymous):

Find the circumference of a circle whose diameter is 10 inches. Use 3.14 for pi.

OpenStudy (gorv):

circumference=2*pi*r =2*3.14*5 as r= diameter/2 =3.14*10 =31.4

OpenStudy (gorv):

@dolphinpink16

OpenStudy (anonymous):

Oh yeah i forgot i had to multiply 3.14 by 10 thanks